OpenStudy (anonymous):

Can someone help really quick?? Find the area of a regular pentagon with apothem 15 meters. Show all the steps to explain it to me please?

OpenStudy (anonymous):



OpenStudy (anonymous):

right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now we have to find the area of:

OpenStudy (anonymous):



OpenStudy (anonymous):

and multiply by 10

OpenStudy (anonymous):

No, the way I have to do it is i have to find the length of a side of the pentagon, and then use area = 1/2ap <--apothem, perimeter

OpenStudy (anonymous):

ok. the angle at the top of the triangle that i drew is 36 right ?

OpenStudy (anonymous):

so the base of the triangle is 15tan36

OpenStudy (anonymous):

so the length of the side of the pentagon is 2 * 15tan36 = 30tan36

OpenStudy (anonymous):

I'm so lost, but go on

OpenStudy (anonymous):

nothing to do more in fact.. we found the length of the side

OpenStudy (anonymous):

but tell me where you lost me and ill try to explain

OpenStudy (anonymous):

So wait, what do I use as the side? i dont just add 30tan36 five times to get the perimeter. would it be like 21. something or what?

OpenStudy (anonymous):

yes.. 5*30tan36 = perimeter

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

ok ? you need explanation ?

OpenStudy (anonymous):

Lol i'm very confused right now.

OpenStudy (anonymous):

why ?

OpenStudy (anonymous):

Did u get my message? :/